OpenStudy (anonymous):

integration for : cos(2x)*cos(3x) please tell me the answer

OpenStudy (anonymous):

I hope it is from -pi to pi

OpenStudy (anonymous):

no its indefinite integration

OpenStudy (jamesj):

\[ \cos 2x . \cos 3x = 1/2 ( \cos 5x + \cos x ) \] Now just integrate that.

OpenStudy (anonymous):

why its not = 1/2(cos 5x - cos x )

OpenStudy (anonymous):

use double and triple angel identity

OpenStudy (jamesj):

cos(x+y) = cos x . cos y - sin x . sin y cos(x-y) = cos x . cos y + sin x . sin y hence cos(x+y) + cos(x-y) = 2 cos x . cos y \[ \implies \cos x . \cos y = \frac{1}{2} ( cos(x+y) + cos(x-y) ) \] Thus \[ \cos 2x . \cos 3x = \frac{1}{2} ( \cos 5x + \cos x ) \] (remember cos(-x) = cos(x) )

OpenStudy (anonymous):

yes i noticed that cos(-x) = cos (x) so cos(-x) will be cos (x) thank you very much james

OpenStudy (anonymous):

ok tell me sin (-x) = ? tan (-x) = ? cot (-x) = ? sec (-x) = ? csc (-x) = ?

OpenStudy (anonymous):

sin (-x) aslo = sin (X) ?!

OpenStudy (jamesj):

sin(-x) = -sin(x). And as you already know cos(x) = cos(-x). From those two identities alone, you can derive all the others you are asking.

OpenStudy (anonymous):

aha thank you very much for your helping and caring

OpenStudy (anonymous):

james are u there ??